OpenStudy (anonymous):

2/5, 9/10, 7/5, 19/10, 12/5 Find the common difference between the terms of the following arithmetic sequence.

OpenStudy (anonymous):

\[1\div5\]

OpenStudy (anonymous):

\[\frac{12}{5}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

Common Difference of an A.P = Second term - First term = Third term - Second term = ... = k

OpenStudy (anonymous):

Here Common difference = (9/10) - (2/5) = 1/2

OpenStudy (anonymous):

\[\frac{12}{5}=\frac{2}{5}+(5-1)d\]\[\frac{12}{5}-\frac{2}{5}=4d\]\[\frac{10}{5}=4d\]\[d=\frac{\frac{10}{5}}{4}\]\[d=\frac{10}{5}*\frac{1}{4}\]\[d=\frac{1}{2}\]